Raven Base Station

(C) Graham Swallow version -1b, gen 1996-07-29 00:37, gps@trix.dircon.co.uk

Base Station Zero - Binder for Issues, copywrite notices, overview

1 Raven Base Station

Space Station Raven is a Linux workstation ( running X11 ). Space Station Raven is a £3 Internet shareware magazine (three UK pounds).

1.1 This is Raven Base Station Zero.

The Base Station is where all the Issues go. Raven Base Station Zero isn’t an issue, it is more like an overview and binder that brings several Issues together. It is the one you don’t have to pay for ¡grin¿, Just as well really, because it is full of such really interesting sections as copywrite, shareware distribution rules, DISCLAIMER, do not spit signs, and access info. It is referenced from all issues, until there is an update! Hence the name. If you are having problems with your browser, get the text formatted version, and read section ”9.3 DOS Readers”. If you are having problems getting access to Raven, you should read this now. IE can you download the files to browse them locally? How can you read Raven without a browser? Otherwise, you are all sorted to go browsing, and can read all this later. Please do make a point of going through it properly ONCE, but since you’re saving that joy for later, all you need to know now is that Raven is shareware, costing £3 per Issue. Enjoy!

2 Overview Conents

2.1 What is Raven

Raven is an on-line shareware £3 magazine to promote Linux, and help you to do something useful with it. Here is a manually edited table of contents. One day it will automatically re-use the same typos as the originals.

2.1.1 Raven Issue-1 You want to set up a Linux box, and marvel at the delights of modern technology. To do so you must already be familiar with unix system administration. To learn that you must have a working unix box. Where do you start?

2.2 Issue-1 - ”Buying and Building a PC” (Intel 386, 486 or more).

• intro - Space Station Raven

• hardware - Plan £1503 for a work station

• hardware - More for Big Budget Video editing

• assemble - Plugging it together

• ip addr - Think of a number, add zero, your IP address is ..

• sysbuild - Boot and Hoot

• fdisk - 1024 sorted

• quick map - list of thinks to do

• part hda4 - /etc/lilo.conf and /hda4 - 10 M below 1024 cylinders

• swap space - Swap Space

• twin boot - Twin Boot Upgrade

• x11 conf - Linux X11 installed and up

Yes - you need to budget £1500 for the workstation. If you have more, get a better machine, but don’t under-estimate other costs. If that is too much for your budget, you should be able to buy a proportion of the components you need for about £800, and get the rest later. You may need assistence from a friend (or work) with a second machine.

2.3 Issue-2 - ”Now what’s this? Linux User’s view”

So you’ve got Linux running, and want a quick start on various things. Such as? Issue-2 has several things that might be of interest to you. Most of them are available on CDROM, so you can do something, if you haven’t got Internet access.

• intro - Issue-Two

• user - User gps

• backspace - BackSpace

• pkg install - xearth - EG Package Installation • sunsite - sunsite: disk3

• sunsite2 - Sunsite Packages

• xloadimage - xloadimage-4.1

• mc - mc-3.2.1

• ddd - ddd - gdb data diaplay

• auis - AUIS - ez - WordProcessor

• lyx - lyx-0.8.4 - Word Processor

• thot - thot

• exts - dvi .ps .tgz .shar .me .ps

• gs - Ghost Script

• gnuplot - gnuplot for botanists.

• archie - archie -via- www

• chinese - Chinese xterm

• kaffe - kaffe - Java Compiler

• cbb - cbb-0.64

• xfinans - xfinans

• pigeon - pigeon - October

• X11 - Linuxing with X11

• urls - URLS to surf

2.4 Issue-3 - ”Internet Working”

Issue 3 gives you a ’drop-in and change the values’ configuration for a connection to the Internet. It documents all the files you need to setup, by providing a ’live’ example of a machine with a modem and an ethernet card.

• efax - fax

• internet - Internet

• modem - Modem

• tcp sys files - TCP/IP system files

• tcp interfaces - loopback • rc inet1 - /etc/rc.d/rc.inet1

• dial ppp - dial ppp

• slip - SLIP

• knl devs - Kernel devices

• tc sys files more - /etc/services and other files

• rhosts - networking with/out passwords

• X11 netw - Networking X11

• tcp apps - TCP/IP apps

• sendmail - Sendmail

• sendmail filters - Configurations on the end host

• knews - knews

• xrn - xrn

• apache - apache www server

• other stuff - Other stuff

2.5 Issue-4 - ”Upgrade to Linux 2.0 ”

You’ve got to upgrade modules, PPP, and of course recompile the kernel.

2.6 Actual Contents

The above list is outline only. For the real table of contents, go back to the actual index.

2.7 File Layout

Although Raven is a magazine, and magazines are ususally ”published once”, and never change. Raven is released as previews, and occasionally updated, so each ”version” has a version number, such as ”-0f”, ”-0g”, ”-1a”, ”-1b”. To allow a site to carry the latest version, and also carry the previous version (in-case I deleted something useful, in the name of progres), all the files have the version number in their direcory path. The site administrator has to point a symbolic link to the ”latest” version, so that you can find them easily, and so that outdated version numbers don’t appear in bookmarked URLS. In practise, the filenames change, and bookmarked URL’s get broken, but I hope you don’t mind that too much. Especially since the ”published” versions appear less frequently.

The first issues of Raven are distributed in the same .tgz file. This is to get you started. Note that postscript files (.dvi actually) may not be in the same file, as they are big. This may change.

./Raven/.
./SSR01/.  --> ssr01-1b # symbolic link
./SSR02/.  --> ssr02-0g # preferred in bookmarks
./SSR03/.  --> ssr03-0g # unless version is vital
./misc/.
eye_view.gip    # created by xearth
gnu_drew.gip    # created by gnuplot + xpaint
./ftp/.                 # individually download-able files
Raven01-1a.tgz  # the out of date .tgz
Raven01-1b.tgz  # the newer .tgz
Raven02-1b.tgz  # the newer .tgz
Raven03-1b.tgz  # the newer .tgz
./ssr01-1a/.            # Issue-01, version "-1a"
./ssr01-1b/.            # Issue-01, version "-1b"
the files for issue 1
./ssr02-1b/.            # Issue-02, version "-1b"
the files for issue 2
./ssr03-1b/.            # Issue-03, version "-1b"
the files for issue 3
ssr03.html      # the index file
ssr03_01_intro.html
ssr03_02_section_two.html
you get the idea, it's a form of invisible version control

3 Overview

3.1 What is Raven

Raven is an on-line shareware £3 magazine to promote Linux, and help you to do something useful with it.

3.2 Raven isn’t Free

Indeed not! I’m hoping to make money from this project. Initially enough to get by without resorting a ”proper job”, but later I’d still like to retire with a pension, thankyou very much.

However, I don’t discriminate WHO is using it, whether for business or at home. If Raven saves a student from 2 wasted days, the value is much the same as anyone else’s time, and the cost is affordable.

Initially, Raven avoids depending on anything that isn’t Free, so it’s value is self contained, and doesn’t accumulate restrictions. I am looking at ways of keeping it within these guidelines, but also developing things of value. These are likely to be ”useful menu utilites”, that you could live without, write yourself or find alternatives from the Net.

In addition, once you have purchased a copy of Raven for (you), it is yours.

3.3 Feedback needed

The chances of getting anything done is remote, but you never know. Send me an email, saying what you think. I might lose it, but if you see that idea implemented, award yourself a warm fuzzy feeling inside.

At present I am torn between leaving Raven as-is, and ”maintaining” Raven, to cover new facilities, and new versions of packages as they happen.

EG a review of Lyx-0.8 is no good to you if Lyx-0.9 is available. However, I do try to put version numbers in there, so you can check for yourself.

You may find that some articles do get updated, and others don’t. Please tell if this is a problem, or if I miss the point somehow.

3.4 Contributions

To keep things simple, I initially want to avoid using contributions from elsewhere (and adverts too!).

The best form of contribution, is to do it yourself, and tell me about it. For example, if you think that my LAN configuration needs a worked example of a multi site sendmail gateway, interlinked with dial-up Internet. Why not write it up, put it on your home WEB pages and nag me to write a paragraph and X-reference.

Then you can select what License your work is released under, but still have it available to Raven readers (who mostly have NET access), and other people too!

This particularly applies to User Groups, who have built up a site, and would like the Table-Of-Contents reviewed and included somewhere. (Remember the off-line CDROM readers too!).

3.5 Target Reader

You tell me who you are, my guess is one of ...

• Someone looking to buy a machine for work and home. You prefer not to be spoon-fed, (build it youself), so you can re-build it knowing ”what’s in there”.

• A young techie in an office network environment.

You persuade your boss to loan you a pair of machines for 4 weeks, and give you 25% of your time to building them. You will demonstrate a WIN.311 pc accessing the Linux disk and printer, Internet and maybe a few other things.

• A young techie’s boss :-)

• a small business setting up a 5 PC network with Internet and FAX facilities. Planning to upgrade system versions every 2 years, or as new services are required.

• a growing business setting up 5 PC networks with Internet and FAX facilities.

Planning to upgrade something every 4 months, this month it’s structured wiring, next month it’s the AUIS upgrade, then it’s back to strategic planning.

• A technician in a large LAN or modem environment, looking to reuse those old 386 boxes as ethernet bridges, routers, printer drivers, modem drivers, X11 terminals -etc-

• A techie on a ”Today: Build a Linux” course.

• A box shifter, with customers asking about Linux.

• A student looking to learn about the Internet, not afraid of computers

• A biology student not looking to learn about computers, but has to.

• A rock band establishing a nymphomation network

• You.

3.6 Your Machine

Probably a Linux machine, but maybe a old Xenix, new FreeBSD, or Apollo, It might just possibly be a 100 VDU SVR4 machine, with several services offloaded to turbo-charged satelite servers. Most likely an Intel based-PC, but not necessarily. Linux has gone 64 bit, and is reported to work on several different CPU architectures. The DEC Alpha jumped the queue, and took Linux on-board with project funding. Contact your local DEC dealer, or one in Linux Journal Here, I’m assuming that you have a limited hardware budget, and can buy new disks, you might also be inheriting an old sun station, or attic PC.

3.7 But I already have a machine

Issue-1 takes you through the motions of building a Penguin from scratch. If you already have a machine, you won’t have to do any of that.

You will probably have to install an extra disk, and maybe upgrade the memory. You can check that that worked, using DOS if you like. You have the CDROM already?

Linux DOES require a lot of disk space to lay out all the apps and play with them, and 2.0 Gig disks cost £200 each these days. £100 for 16 M (making 20 or 24 if your MB can pick-n-mix) £100 modem, £120 Internet £120 phone £100 spinnach £100 books £50 cables, means that Uncle Bob’s ”free gift” of a PC costs £1000 in the first year. Almost as much as half of the rent!

Now follow Issue-1 to install Linux onto that new £200 disk, probably using LOADLIN.EXE until LILO or GRUB works.

3.8 Your Machine

You should have a large machine to see it all running properly first, and to assist the smaller machine (network CDROM).

My ’typical hardware configuration’ is what USED to be called a big PC, but is now quite affordable, and fits in one small box.

tower + monitor + keyboard + wires

If you only have a 386/sx, 4Mb RAM and 40 MB disk, you can still proceed, but life won’t be easy. If you have it as a second machine - you will see a lot more of the unix world.

Rebuilding small machines and putting them to good use, is part of the fun of Linux - which supports a large range of hardware devices.

I have heard USENET reports of 386 4 MB boxes, with no hard disk, booting from floppy and working as an IP router. (YMMV).

4 Online Availability

4.1 Download

If you are reading this on-line over the Internet, you may prefer to download the archive, uncompress it and read it off-line. Raven is intended to be download-able, so that you can fetch it.

It is also intended for inclusion on Freely distributable CDROM’s (shareware). I also have text versions of the first issues, for access whilst your HTML browser is down.

The file that you download will probably be a tar archive file (unless I ate my todo list, and made an RPM version). To expand that into seperate files use the tar -zxf Raven-9607.tgz command, or follow the README.

4.1.1 Version ”-0g” previews

If you have troubles downloading the files, please try different sites, and nag me with email if something stays broken. Your best bet used to be to get lucky and find trix online, now thanks to lth.se, you can browse an over-the-shoulder preview.

4.1.2 Version ”-1a” release

So  far  the  ”-1a”  release  has  been  available  as

a .tgz on the InfoMagic LDR Apr 96 (on disk#6), from http://beyond.mamlo.lth.se/˜ gps/Raven/ftp and ”open-for-reading” on http://sunsite.unc.edu/pub/Linux/docs

Whilst  waiting  for  the  ”-1c”  release,  try  browsing  a  recent

snapshot at http://beyond.malmo.lth.se/˜ gps/Raven/SSR01/Raven01.html . If you have a browsing web robot, use it to grab a local cache copy for offline brosing.

4.1.3 Edit the URL

If you are using lynx, there is no editable URL line, except the bash command line. If you are using netscape, it is easy to edit the URL line (HINT: click on it) and attempt to browse the directory table of contents (some site aren’t browable, and require an index.html file).

If you get a bookmark from a friend, but the filename has changed, simply reduce the URL and try again.

4.1.4 keep your own netscape cache

I don’t have a web robot, but I found that running netscape, on local file://tmp/mnt/phoenix/ over user fs/ftp fs, gave me the latest browser, that kept local copies, even after the phone went down. To do that you need userfs. If you don’t have netscape or lynx, use mc.

Regretably, userfs ftp only works over ftp, and my ”over the shoulder” previews are on an http site. Oh well, either write a robot, or become one and save every page!

4.1.5 mc-3.2.1 as an ftp browser

mc has its own ftp fs, (only works on ftp://site not http://site), F3 can view most types of files, F4 fetches a local copy for you to save, and F8 copies highlighed files and entire sub-directories.

4.2 Where to Find it

4.2.1 Offline at trix ...

http://trix.dircon.co.uk/Raven is my home machine, connected by a modem. It does NFS, SMB, FTP, HTTP and you are welcome, but it does have a limited bandwidth, and is prone to going offline suddenly.

4.2.2 Online at sunsite.unc.edu

http://sunsite.unc.edu/pub/Linux/docs/Raven is the ”published” version. There should be a .tgz file in the Raven/ftp directory, but check the README files, as most people DONT want the postscript files, so they should be seperate.

4.2.3 sunsite worldwide

Sunsite is mirrored in many places http://src.doc.ic.ac.uk/packages/Linux is one.

Since sunsite is such a common service, I see no need to maintain other sites, but if your country or region needs better access, and you wish to provide it, simply send me an email, and we’ll arrange something.

4.2.4 Online at beyond.malmo.lth.se

http://beyond.malmo.lth.se/˜ gps/Raven is where you will find the latest over-the-shoulder pre-publication snapshot of Raven.

You should NOT mirror it, just nag me to release it properly!

You may take a personal local cache copy to browse offline.

4.2.5 CDROM Access

If this is on a CD, you should find a directory of uncompressed files, (ie a mirror of sunsite) and you can browse them directly.

DOS PC’s may have problems with the long unix filenames, when not using Rockridge extensions. In that case view the generated text file, build a Linux box, and be happy.

If you don’t have X11 running, try lynx or the text formatted versions. Issue-One has a useful overview on X11 configuration.

4.2.6 BBS access

When the ”-1c” version is out, I shall put a DOS-ified version on the BBS circuit. I don’t know much about either, so be patient. This will enable people to download something to get started.

4.3 majordomo info list

To get announcements of updates to Raven, you should join the raven mailing list. This is low volume, and (currently) for announcements only. To do so run the following command from a shell,

        echo subscribe raven | sendmail majordomo@trix.dircon.co.uk

4.4 Try Before You Buy

This is £3 shareware, I hope you like it. I expect you to download it and read it before deciding!

4.4.1 Issues with themes.

Some issues have a theme, which makes it easier to select which issues you find relevent to you. So if you already have the hardware with Linux installed, you would probably skim and forget issue one.

Later issues may have a combination of themes. Hopefully it will still be clear to you, whether you want each issue or not.

Remember this is try-before-you-buy shareware, the only warrenty is that you decided it was worth the cover price!

5 Copywrite

5.1 Copywrite

All material in ”Space Station Raven” is copywrite of the author(s), who expect payment.

5.2 Shareware

The selected distribution medium is a shareware magazine, so you can down-load it, read it and use it - before deciding if you want to buy it.

5.2.1 Payment £3 (GBP Three) No discounts.

Payment methods are limited, these are early days.

Hopefully any number of distributors will offer payment facilities to their customers.

5.2.2 Who to pay

        Graham Swallow
        87 Burley Road
        Bransgore
        BH23 8BA
        United Kingdom

5.2.3 What you get

After payment do not expect anything else.

You get the right to print the text (if you want to), and a ’user license’.

5.2.4 Keep Records

Keep a record of all payments, and which Issue you paid for. Every new issue that you enjoy, update the list.

5.2.5 What you Keep

If an issue gets updated, and you have purchased that Issue, the updated version is already yours. (Please Note: it is up to you to decide whether the newer version is better or worse than the version you chose. Don’t simply delete the old!) There is no implied ’updated’ version in the pipeline.

5.3 PRICE - No discounts

The £3 price is already discounted. Any individual gets the benefit of this bulk sale (none sold yet!). User groups may support their members by making it easy for them to obtain Raven sooner, in better shape. If further discount is appropriate, it will come in the form of content.

5.4 PRICE + costs (International transfers)

In addition to the PRICE you must pay any extra costs (such as printing, import duties and money transfers). International transfers are unbelievably expensive at £10 to £15 pounds each. You can arrange payment, but you have to pay the cost of doing so! I can accept UK cheques to ”Graham Swallow”, to the address above. If you have friends in the UK you should get in touch. It is not a good idea to post cash, or even travellers cheques, even by registered post. You are responsible for any local import duty or taxation. In the UK imported goods VAT tax gets collected by the postman.

5.5 Getting a receipt

If you require an invoice (either by post or by fax) that is an extra £2 handling charge (admin overheads-P+P). Any number of items per invoice. However, if you have otherwise paid and wish to confirm safe delivery, and your registration, you can use email (free). Receipt is after payment, which is upto 28 days after intent of purchase.

5.6 Registering your copy

I want every copy sold, registered (unless it is too much work for me). This will confirm that payment has arrived, and may make you eligable for a random reader prize draw (”the 1000’th issue” or whatever).

You should see this as a VOTE of support for Raven, and there may be fringe benefits, such as ”Business Card” database, or name of re-seller.

NOTE: you will (by default) remain anonymous. There will NOT be a mailing list, either email or paper, though this will be a risk if you choose to ”unmask”. Note that posting to a news group, is also a way of making your name, email address and general interests known.

NOTE: if at some stage (not quite yet), you receive a ”Voucher” for an issue of Raven, eg as part of a larger package with that particular issue bundled.

THEN: if you have already purchased that Issue of Raven, (and registered it), you will be able to transfer that voucher to a different issue, or give it away to a friend.

Sorry, I have no facilities for refunds, if this is a problem you must make arrangements BEFORE (not) paying.

5.7 Registration Details

It helps if you can send these by email, (note that US-UK post takes about 2 weeks, email about 2 minutes).

I want your postal town, for a survey of WHERE the issues are going, and I need some unique identifying secret to validate your email. I need:

        Your Full Name
        Your Email address (if you have one)
        Your Postal Address
        Your bothers birthday (or other Q+A password)

If this is sucessful, maybe I’ll send YOU a postcard!

5.8 Site License

There is no such thing as a site license for Raven, but you may wish to purchase 20, and let people in the office, sign up for them, either as their job role, or as private individuals.

If one technician uses Raven to build 100 workstations, you only have to pay for the one reader.

You might like to purchase other issues for the users of the workstations, again either as their job-role, -or-purchased for private individuals. Ask your people what they want, and make it possible.

5.9 Re-selling Raven

I want to encourage resellers to reduce my admin. overhead, but you must apply (by email) to sell Raven.

This is a simple email request, to establish who you are, and keep track of every issue and supplier. There is no charge, this is intended to protect the customer.

Linux User Groups can make a collective action, and purchase several items in one go, but individuals should then register their copy, this can be done by email (preferably by the LUG).

Distributers may re-sell Raven at profit, but there are no discounts for volume.

5.10 Re-Distributing Raven

If Raven is being distributed, and not being sold, it should be clear that this is so.

This is implicit to most anonymous FTP servers, and totally free CDROMS, but distributions with royalty based add-ons should make it clear that they do or don’t include payment for Raven.

If your CDROM is being distributed within your company, or university, you should make it clear which Issue’s of Raven have been purchased, and which have been re-distributed. This particularly applies, if you are pointing readers to follow sub-sections of Raven, when the reader should be reminded that it is (un-)registered shareware.

5.11 License to print Raven

If you sell a customer an issue of Raven, the customer gets the right to print it.

You may assist the customer, providing a double sided bound copy. You should not change Raven, or add your logo except for Printed By:

This license is granted to any shop who print less than 100 in 1996, and expires at the end of 1996. If you print more, or after that date you must request permission to do so. This is to protect the customer.

(You must still apply to sell Raven).

Question: If Raven was printed in the LDP compendiums, would you then be prepared to pay for it? Or is that idea unacceptable? (Flames to c.o.l.misc or advocacy).

5.12 Every issue, is release on an ’as-is’ basis.

You read it, you decide.

5.13 No Future Issues - ? - (Delete this section)

I hope to develop Raven, with new issues, but who knows. Although Raven refers to itself as a ”magazine” or ezine, there are no definite plans.

If someone offers you a ”subscription” to future issues, report them to the police. If you send money in advance, there is no repayment. This is shareware:- try BEFORE you buy.

The basic proplem is that if I go back to contracting, there won’t be any time for future work, as well as a decent life. Vote with your dosh.

6 Copywrite and DISCLAIMER are in the Base Station

The Copywrite note and DISCLAIMER notices are currently common to each Issue, and appear in the Base Station. Click above to find them.) You can locate Raven at

• http://sunsite.unc.edu/pub/Linux/docs/Raven/

• http://beyond.malmo.lth.se/˜ gps/Raven

• http://trix.dircon.co.uk/Raven

• http://www.users.dircon.co.uk/˜ trix If you are having problems locating them, please contact me by email. gps@trix.dircon.co.uk Raven is copywrite Graham Swallow, it is try-before-you-buy shareware. The authors cannot accept any liability for the content of this magazine. It is up to you to check that any actions you take, makes sense, matches other documentation, and work on a test system. This is a summary. Please locate and read the actual notices.

7 Free Distribution

Raven can be distributed as shareware.

7.1 Mirror Sunsite (current ”published” home)

You can find Raven on sunsite.unc.edu in the docs directory. You may mirror it verbartim, or individual files from the ./Raven/ftp/. directory. If you mirror sunsite en-masse, do so! No problem. You must remain up-to-date, and after 1 year you MUST either delete the files or move them to ./old-Raven/. The package must remain intact, you can un-tar all the files for use, you can export (all) the files over ftp or www, you must not modify them. Note that the packages are modular by Issue, so that’s easy. You must make it clear that this is shareware, and that you don’t (or do!) arrange payment for it (implicit to most ftp servers). (A sample cover page link might help here?).

7.2 put Raven on your BBS or CD

You can make Raven Issues publicly available on the Internet or other BBS systems. It may be included on CD’s. Please notify me by email, and only distribute recent ”official” versions (not the previews).

7.3 put Raven on your NFS, FTP or WEB server

If you wish to offer your users access to the HTML pages, you may do so by installing to a directory like:

• http://localhost/www/Raven/ - Space Station Raven - A Linux workstation

• Raven/ftp - Raven is £3 shareware, download it

• Raven/ftp - Purchase it here - £5

• - Printed 70 pp £12

• - web 200 K

• - dvi 600 K

• - ftp 1.6 M

7.4 put Raven on BBS nodes, or floppy disk sets

I think you get the picture, if you re-distibute for Free, there is no charge for doing so, however you MUST STILL APPLY TO DO SO.

When I understand a bit more (than nothing) about the FIDO-NET, ie when I can automatically update, or withdraw a distribution, I will do so, but until then simply ”apply” by email to carry Raven.

NOTE the following about commercial BBS’s:

7.5 Commercial BBS’s – VITAL NOTE – !!!

If you offer your customers the service of paying the registration fee of any package, you must offer the SAME SERVICE for Raven, even a single issue.

You may charge a fee, and make a profit, and add that to the price the customer pays.

7.6 Commercial BBS’s – VITAL NOTE – !!!

If you use a toll-call system, (in the UK that’s an 0891 number), or if you typically sell (other) packages for more than the raw price of a single Issue, you must take responsibility for collecting the cover price, and forwarding it to me, along with sufficient details about the customer, that they could later claim that purchase.

For ”typical package distribution cost” take a simple site average, and include any membership fee, or BBS subscription fees.

If you do so, you may advertise that the cost already includes the full registration.

This appears to go against the tide of try-before-you-buy, but hopefully, there will be few un-happy customers, and those who require their money back (upto 28 days after purchase), get it.

They should send their request to either/both of us, preferably with a brief note why (for QC).

7.7 Commercial BBS’s – VITAL VOTE – !!!

Don’t get me wrong, I have nothing against Commercial BBS’s, and shareware is central to BBS’s, and this is shareware, but even so the try-before-you-buy rule doesn’t apply! You can make this clear to the Customer. Instead the money-back-guarantee applies, with bundled convenience of pre-paid purchases, and peace of mind (one stop shop). The reason for this is the low unit cost, and the other general availability of Raven (eg CDROMS).

7.8 Commercial BBS’s – VITAL NOTE – !!!

If you think your ”typical package cost” falls below the threshold, and do not want the administrative overhead that this involves, but you are uncertain, please send me a sample price list by email, and get clarification.

7.9 NO Printing!

You may NOT print out the magazine unless you have paid for it.

8 DISCLAIMERS - NO WARRANTIES

Please note that in the UK, our infant reading schemes do not use capital letters, but word shapes (flashcards). This means that proper text is more readable than everything in UPPER CASE. Personally I cannot read blocks of text that are all in CAPS LOCK, a mild form of dyslexia. I can’t read it and won’t write it. I am relying on your intelligence to perceive that this is the DISLCLAIMERS section, so it is not converted to all upper case. Similarly it should be clear that this applies to each Issue (this is the Raven Base Station). IE A disclaimer would only make sense to an adult, and an adult could easily find this section, and read it.

8.1 WARNING

8.1.1 Risks There are risks such as losing data, losing hardware, and there is a risk of injury or even electrocution.

8.1.2 NOT Mine The authors and publishers of Raven cannot be held in any way responsible for your actions after reading it.

8.1.3 Yours This magazine is for people who want to build their own workstation from scratch. There are risks, and they’re all yours.

8.1.4 Choices and Rumours

Recommendations made in Raven MUST be taken as sweeping generalisations, for you to think out. Remember, you are the system administrator - not me.

8.1.5 Test it to your OWN satisfaction

You can minimise the risks and costs, by not depending on a system that you have not tested to YOUR satisfaction. You build up a new system to your own specification, test it, then depend on it according to your own judgment. When happy, you bring the new system into use.

Be prepared to start again from scratch, or not be able to do everything you used to. With the old system to depend on, you have only lost time and progress.

8.1.6 SIMON SAYS

To allow a more ’direct’ style, which is less verbose and more likely to be remembered and understood, Raven will tell you what to do in the tersest way. You are expected to question every instruction, and make changes to suit your system.

Above all you should locate the proper documentation, and use that in preference to Raven. Hopefully there will be only mild confusion.

unix systems is a complex subject, and easy to get wrong. Any errors in Raven are errors. You should not use suggestions taken from Raven before testing them on a test system, and making your own mind up. See the unix.horrors faq.

8.2 Phew!

Nuff said.

8.3 Postal Address

        Graham Swallow
        87 Burley Road
        Bransgore
        BH23  8BA

9 Built with Linux (DOS readers)

9.1 Built with Linux Doc SGML!

Raven is produced using freely available tools running on Linux.

For more information on SGML production, see a later issue! For those who want to know more now, look in sunsite /cdrom/utils/text/linuxdoc-sgml-1.5.tar.gz

linuxdoc-sgml takes a single file.sgml, and produces one of several output formats, for sending to a postscript printer, use on the web, or even a line printer.

Each of the formats is neat (well you judge for yourself).

The .txt file can be read by the less utility, which interprets backspace (”reverse motion control chars”) to give bold. These simple .txt files are supprisingly easy to read, and navigable with a printed table of contents.

Linuxdoc-sgml also generates the .html pages, though the file names have been renamed. It generated the default table of contents top page, with one html page file per section. By default the sub-sections are numbered 1.2, and appear in the top index.

linux-doc sgml also produced a TeX source file, which it delivers as a pre-compiled .dvi file. That can be viewed by xdvi, and it looks excellent, but html is a lot faster.

Some people have established their own TeX profiles, and macros that they trigger with SGML tags defined in a customised DTD. However, that looks a bit complicated, and sometimes it’s better to stick with the ”supported” established default configuration, and treat it as a ”predictable sub-system” that gives you consistent GI-GO every time.

You then build a convery belt of filter/converters such as:

        generate .sgml source through M4 and sed  filters,
        generate TeX.dvi output using patched mapping files,
        rename generated .html files from .toc

The technical document layout is a tad boring for a magazine, but helps to make the text have a consistent style. If you write ”numbered paragraphs” style documents, it is worth investigating, but be aware that you have to adopt it’s overall indexing style, rather than adapting it to you.

9.1.1 Browing Authors

If you produce your own WEB pages, you may have some opinion on the production. To properly show off Linux capabilities, I should be using a web oriented page editor, and producing dramatic looking documents.

If you use a really neat production system, gimme a call, I suspect that somewhere in the ’tk’ archives, is a rock stable word processor and html page layout designer. I haven’t found them yet.

For the time being, I’m happy with linuxdoc (and if not - why not). As for the graphics content, the eye view .gif was generated by raven using xearth, then xpaint added the title lettering.

9.2 mkdir Raven

After down-loading the .tgz file(s), untar it and browse the files. This will create a directory called Raven and the files in it.

        cd $HOME
        tar -zxf Raven01-1b.tgz
        mosaic Raven &
#
#       you should now be browsing Raven Locally
#

9.3 DOS Readers

If you are a DOS user, you will want GZIP.EXE, GUNZIP.EXE and/or GTAR.EXE, or other UNIX-DOS utilities. But you knew that already.

If it works, use a browser, such as Mosaic to view the html files. If the filenames are all mangled, or it just doesn’t work, read the .txt file (I have :-). It’s a but monolithic, but it works.

The .txt file looks strange with some browsers, and crystal clear on others. If you see BBoolldd TTeexxtt (Bold Text), and think it’s ugly, look again with a different tool. If you have Linux running, use less.

Remember to locate your resources, does your browser work ?

9.4 mc # F3 filename.txt

mc can understand the backspace formatting, but you might have to tell it to try, using the F9 key to (un-)format the text.

mc can also browse compressed man pages, and wander around inside compressed tar archives.

9.5 less filename.txt

If you don’t have a browser running, the files are available as mildly formatted text files. The Linux utility less is very good at displaying backspace formatted text, you might find something that works under DOS, or get rid of the formatting.

9.6 ddoouubbllee vviissiioonn – bˆ Hboˆ Holˆ Hldˆ Hd

It looks good when viewed with the less or zless utiltity, but dreadful when looked at in an editor, or most HTML browsers.

Bold Text is done as Letter BackSpace Letter, so that on plain printers, (dot matrix), each letter gets printed twice, and appears bold. Within editors, you also see the ˆ H character making it almost unreadable.

9.7 getting rid of the backspace formatting

You can convert the files to boringly plain text files, using using sed or vi.

These may also be available on DOS - CICA or elsewhere.

        sed 's/.^H//g'  # substitute /Any-Letter BackSpace/nothing/g-flag

If you have trouble typing CTRL-H to the command line, prefix it with CTRL-V to tell bash that the next character (key) is to be taken as data.

If typing this command is difficult, put it into a script and run it. Remember to tell it to read from the input file and pipe it to less. You can also use a supprisingly similar syntax with vi

        : 1,$  s/.^H//g

The quotes are optional (elvis allows them), and you still need the CTRL-V key. If you put

        map #1 : 1,$  s/.^H//g

into your $HOME/.exrc it might make the F1 key do it (if it isn’t already busy).

10 What is Linux

10.1 What is Linux

Linux is not U**X. Linux is a POSIX based kernel, that makes any decent chip, starting with Intex 386, look like an operating system.

When combined with the many other components, such as libc, init, grep, gcc, tcp/ip, sendmail, (etc); Linux looks like a unix operating system, with loads of development tools.

When combined with the standard free X11 toolkit, and several substantial and sucessful free applications, like tcl/tk, /usr/andrew, ghostscript, TeX, it looks like a workstation.

Linux has been ported to other CPU’s and architectures other than Intel, and you should check the state of progress of your favourite machine. Raven assumes you have a 486 ”AT”, because it’s easier that way. Remember that almost everything is available as source.

If you don’t have Linux, but do have FreeBSD or SVR4, most of Raven is still relevent to you.

Linux is written by and Copywrite of Linus Torvalds, available under the GPL, but with caveats about it’s intended usage that make it almost Library-GPL (but not). Many people have added their driver code, which are ususally bundled with Linux (for convenience of fetching, building and maintaining them at a specific version).

Linux 2.0 is a recent release (June 96). It is a progression of 1.3.90, and substantially better then 1.2.13 (eg mmap files).

10.2 But Linux is only the kernel

This is true. Linux is just the /usr/src/linux kernel, that provides really interesting features like fd = open(filename, mode), virtual memory, device drivers, filesystems and process management.

There are many other packages that have been bundled under the Linux banner, that are required to help make Linux look like a unix system, or make Linux possible. These include utilities like ”grep”, system libraries like ”libc”, and essential programs lime ”/etc/init”.

Traditionally, unix systems are released as, and named after their kernel release. Hence Linux.

10.3 Is it really unix?

Not really, because that is a trade name and product. Linux is a clone of UNIX, following many of the various standards.

But it looks like unix to me.

You only have to look at the rate of its progres to realise that it is neck-and-neck with real unix, and moving very quickly. It might even become a de-facto unix, that others have to match. Probably by sheer weight of numbers. Of course, the existing unix standards are stable, so don’t expect Linux to change them by much.

NOTE: Most people use UNIX to mean the official SVR4 product and the trademark, and Unix to mean the generic familly of Unix-like systems. Some people use unix to mean the generic familly, as most trade marks are names and have an initial upper case letter.

10.4 So it looks like unix?

Yes, and from several angles.

Programmers can write code that runs on Linux, that also runs on other unix varients, and clones, with few changes. Look inside a few Makefiles, and you see the different targets that that code supports.

More-over, programmers take their work home, and make it compile on Linux (or bring Linux to work, and make a home from away from home).

System administrators will often see the same utilities that run on SVR4 machines, doing the same tasks. eg when the system boots, fsck scans the disks, just like any other unix.

There are a few major libraries that aren’t supported on Linux, but then they aren’t supported on many other unix’s, and programmers still code for the alternatives (usually).

10.5 unix - more or less

SVR4 has a kernel level streams architecture, that can be very efficient and flexible. Linux doesn’t have that, but Streams in itself can be a sophisticated protocol, with many unused test cases that have to be programed for, to be complient with the standards.

SVR4 (and SVR3) has the ”TLI” transport layer interface, that is supposed to abstract away from TCP (whilst running TCP/IP underneath). Linux doesn’t have this, just using TCP/IP directly. However, many programmers would question the absolute value of TLI, given that it already has an installed base of required ”features” and workarounds.

SVR4 has a few ”TP” transaction processing libraries, that Linux doesn’t have. These are sophisticated database back-ends, that are like a mini-operating system, knitted onto the core operating system.

Linux has a few features that SVR4 doesn’t have. In particular loadable modules, are kernel drivers, that can be compiled and loaded without recompiling the kernel. Proprietry device drivers can be issued on a diskette, and installed without re-linking the kernel, taking about a second to load.

Linux networking has some features, such as tunneling, that other OS’s could only dream of (if only the budget were available, and customer demand, ... ).

Linux filesystems are more openly documented than SVR4 VFS (by example mainly), and Linux has user fs, where the fs can be implemented by a user level program, not a kernel module.

10.6 Development tools ?

The original killer app, is GCC, the GNU-C-Compiler. It is also a C++ compiler.

The argument for Commercial based development, falls down miserably when looking at the plain K+R C compilers that just wouldn’t go away. Of course people still check that their code still compiles and works with other reference compilers, but GCC is everywhere.

GCC combined with the libraries, and standard kernel API, makes it easier to port utilities to Linux, simply by compiling them. And when the current round of testing with 64 bit addresses is done (32 bit has been common for a long time), those tools and their code will be even more portable (with even more #ifdefs in them!).

Other languages tend to be written in C (compilers), and arrive as soon as GCC can make space for them. TCL/TK for example is delivered as source, and compiles on Linux, and SunOS, and ... Suddenly you have a scripted GUI, (running tcl scripts), that you can incrementally edit and run, without exiting the application under test.

Other languages include perl, ada, eiffel, fortran, cobol, lisp, and all the libraries that have been developed for them. Either free or Commercial. There are also XBASE (from DOS dbase-III) compilers, and SQL interpreters.

As well as compilers, you (may) need libraries. Motif is commercially available, Tcl is downloadable over the net, as are many others. Basically, if your in-house system needs a library that isn’t available on Linux, preferably with source, then you should be asking what would happen in 5 years, when you need to expand your system throughput, but don’t have many options open to you.

10.7 Is it as good as DOS

Not quite yet. The problem at present is the number of Office Applications, and finding ones that your people are familiar with, that run well on X11, at a comparable price.

For some, it is already better, because they have learned to use it, and can do everything they want to do. They can produce TeX documents, containing gnuplot diagrams, produce WEB pages, etc etc.

It’s really a question of what your application is. If you have an in-house specialised system, that is a generalised package for your trade, you need your software house to come up with a version for UNIX.

Linux can also be quite bewildering. A reasonable unix installation, requires as many directories as DOS needs files!

10.8 Source of course!

Almost everything on Linux has the source code available, either to the general public (eg Qt) or to software houses with the money. ”So what?”, you say, ”I never read it”.

You are already living within the limitations of your existing software package, so any bugs are likely to be ”acceptable”. Provided you can accept staying within the same hardware and OS base. But what if you really want to get something fixed, or identify the fault?

It won’t be cheap, but if you have the source, you can pay a software house to fix your problem(s). If you are running a dentists business on basic-82 (fiction), you can either rewrite the (source) application to port it to run on basic-97, or you can fix the problem in the basic-82 interpreter. If you are a chain of dentists (eg 20 independent sites), you can form a user group.

Even getting the specification of your existing data, may be enough to migrate most of it to a new application, running from scratch.

10.9 So it’s as good as NT?

More or less, yes.

NT has it’s own toolset, that differs from linux. If you use a compiled D-BASE-III application, chances are you can find a matchable engine to port your application to, on both systems.

Maybe not if you are using lots of NT windowing tools. Then you are well trapped down that avenue, and should look at converging, eg via GUI libraries, SQL common subsets, or your own C++ code.

If your service is completely independent of the OS, such as being a WEB server, Linux is already a better option, with Apache, Sendmail, FTP, SMB, NFS, ... You can transfer the WEB server functions from an NT WEB server, onto a Linux WEB server, and not notice the difference, except that both services ran faster (two CPU’s etc).

However, if you start from no installed base, and have to write code either for unix, or for NT, you might be better writing it for X11, and putting an X11 (workstation) interface on the NT.

If you wrote it in TCL/TK, you could make a few diffs to make it run on Linux-X11 or WIN-311, from the same source.

10.10 What about Networking?

Linux, uses TCP/IP over ethernet (or slip, or ppp), so it is like most of the non-dos world. It can do novell (recent), it can talk to macs, it can client or serve WIN.311 shares (disk, printer, process) over TCP/IP, but not over netbios.

Linux TCP/IP is a full-feature system, and it’s growth in numbers will get it tested in all sort of circumstances.

Being source, the serious hackers have devised various tricks to make TCP/IP work in new ways, such as IP Masquerading, tunneling, etc.

Linux makes an excellent WEB server, it has a 9% presence in the world population of web servers. That is more than NT and less than Sun. Development is on-going, which means that you can run Linux-2.0 until Linux-3.0 is ready and tested (tested by you).

10.11 Linux networking interfaces

Linux supports many different ethernet cards, and with the pending huge numbers of home dial-ins, it will get some testing against the standard PPP servers.

ISDN devices are more expensive, and tested more selectively, but more cards have yet to be designed, and Linux stands as good a chance as NT or SVR4 to get supported.

Fibre links, and expensive lines, tend to come with their own expensive routers, requiring less interfacing to connect to the Linux box, or maybe is Linux-Inside!

10.12 Linux PCI Card Interfaces

Linux already supports many, many PC cards. New SCSI cards and other cards are appearing every month. The companies that sell them typically put the DOS drivers first, but the gap is closing, and becoming more obvious when it is there.

You still have to check for compatability, and reported reliability, but if that SCSI card works for you, then it works! If you don’t change anything, it will still work tomorrow, and if you have to change SCSI controller completely, there are alternatives.

10.13 Multi-Media Internet WorkStation

It’s really a question of finding apps that you can live with, and getting used to using them.

You can send and receive email. you can write robots to do that for you.

You can browse email, news and web pages, switching between them by clicking on URL’s, and buttons.

Finding Internet tools that really suit you, can take time, but Raven (Issue-3) has a few suggestions, and you can start looking immediately.

10.14 Multi-Media Text Processing WorkStation

This means running something like TeX or roff, and the surrounding macro libraries and fonts. These convert plain text files, with embedded text commands to typeset page layouts, with kerning, characters, tables and simple drawings.

The output from either (Tex, groff), tends to be .dvi or postscript (though groff is still very good at formatting to VDU text formats, as used by the man pages.

The input to either, is one of (1) manually edited, (2) edited using a wysiwyg editor or (3) a pipeline output from a previous generating phase.

In addition, the output may be HTML text files, so these utilities are starting to act as filters, or macro processors (which is what they are).

10.15 Multi-Media Graphics WorkStation

Browsing a .dvi file, onscreen is a form of graphics, as is converting a .ps file to an X11 pbm bitmap.

Displaying a static graphic file, can be done by converting to a nearer format, and using the native display library to do it.

Graphics workbenches can can process scanned photos or generated diagrams, (from programs like gnuplot, or tcl/tk). You can then include them as attachments to messages, within printed TeX documents, or within .html pages.

3D grahics programs can take ”standard texture descriptions”, and wrap them around ”standard triangulated surfaces” to render objects under lighting, or use Ray-Tracing to POV ray scenes.

There are standards, at various stages, and Linux has Free Libraries that can do many things. Linux differs from DOS, because in DOS you tend to get one application that does everything, with Linux you get lots of adaptors, converters and utilities. By using them together, you get a tooled-up workbench, though at times it looks like a jig-saw puzzle.

10.16 And this is Free ?

”Free” is of course a subjective term, perceived as ”no restrictions within the horizon”.

There are several ”Free” licenses. GNU popularised GPL and Libray-GPL, (which you get the chance to read several times). Many authors have adopted the GPL, as being an established umbrella, to License their code for ”Free” use.

A lot of X11 software is Contributed by large companies. It isn’t GPL, but is is Free.

Some software is Released, but ”not for commercial use”. This is particularly troublesome, as a typical computer (wordprocessor), used for a home office, is arguably being ”commercially used”. Some Licences clarify that they mean ”commercially sold”, but you have to read them to be sure.

The GPL itself, slows down the free re-use of software, by making it difficult to re-use useful portions of code, in other packages. It could even be argued that GPL code, cannot be used in Library-GPL packages, except where ”bundled” as independent objects. However, the benefits from the GPL, are already obvious, with so much software using it as a license.

Personally, if Free is so hot, why are any restrictions needed at all? Public domain means that it belongs to everybody and anybody, almost without restriction, but in this cynical world, many people think that Free -plus- marketing isn’t enough, restrictions are needed to protect ¡something¿.

10.17 Check your freedoms

With current legislation, you have to make reasonable attempt to be sure that you have Licenses for all your components.

If you are getting your software from CDROM’s, the distributor has usually checked that every package is freely distributable, but they may only be watching their own back, as they are doing the distribution. You may have to comply with every package’s USER License, as well as the distribution rules.

This particularly applies to CDROM’s that come with a few single user Licenses. Official RedHat comes with a Commercial-X server. Others come with proprietry databases, or desktops, or combinations of packages, bundled into an affordable CD.

Some packages have the opposite sense, they are freely usable, but not distributable! They appear on FTP servers, and you download them yourself.

11 Why Slackware?

11.1 Why Slackware ?

What is the best distribution? Why is Raven based around Slackware, not Red Hat or FT or debian ?

No reason at all. I just happened that way.

My first Linux was a SLS set of floppies, I loaded them, played with them, and did other more interesting things. Eventually I decided to upgrade my system to a CDROM, and see if any real apps made it worth double-geeking my life.

I dumped the diskettes on a friend and got a Yddgrasil CD, a CDROM, a modem and (err ...) a phone line and Internet account. I resigned.

The Yggdrasil came as a CDROM and a boot floppy, and they both worked, with ZERO installation. The hard disk wasn’t even needed, until my 8M and X11 reminded me to create a swap disk. Everything ran from the CDROM. To get the extra apps I subscribed to InfoMagic’s Linux Developers Resource, (it certainly has been), and saved a lot of time on the modem.

As a gesture to work, I demonstrated a triple boot trix, with Internet and ethernet. It had WIN311 in 150 M, SVR4 in 280 M and Linux on an 80M Syquest cartridge (I took out the system and handed it to them). It had all TCP apps working, but regrettably I only had the screen at 640x480 (512k Gas Plasma display). Whilst tooling up I got a second machine (raven), and stopped having hassles when cold building trix.

Eventually, the postage system got my resignation through to my boss, so I started on the Raven project in ernest. Then the house became un-inhabitable, rescheduling my sanity as well as Raven. A twisty turny path later, I’m still using Slackware (and still looking for somewhere to live).

The Slackware upgrade to ELF (long ago), was the simplest most natural step, and way ahead of the opposition. I was impressed (though I did a build from scratch).

InfoMagic don’t produce distributions, they copy them to CDROM, and distribute them. The more recent upgrade from Kernel 1.2.13 to 2.0 (eg via 1.3) was not particularly well supported in any release on the Apr 96 snapshots, but when Linux 2.0 was announced, the others soon announced their new products.

Infomagic LDR doesn’t come with support, but InfoMagic had already fixed the bug I sent them (seems fair, they sent it to me). And they sent that response quickly.

Other than that Slackware works.

12 Costs

12.1 Why isn’t Raven printed?

I did look at the cost of photocopies, but got quotations of 20p per double page! I did a quick calculation for printing them on a personal laser printer, but even then the costs were very high, and the additional cost of employing someone to make the copies was unrealistic, as well as the huge amount of time and disruption that it would occupy.

Hopefully you can read this on your workstation, but I do realise that people buiding their box might not be happy with that.

Box-shifters out there, wishing to offer their customers the facility of accessing Raven on paper, should contact me by email.

Printed copies may be available in future, but at present I don’t have the facilities. I’d prefer not to cut trees, or generate plastic wrappings for incinerators, but these are options, especially when volumes will make this more efficient. However, that may be some time.

12.2 Lets keep Raven unprinted

I know that printed documentation is MUCH easier to read, but lets avoid doing so, if it is practical.

However, this is not a religious war! Things like tables of contents, annotation sheets and reference cards, should be printed to get the most from them.

NOTE: the tables of contents can be printed by netscape to postscript files, and printed as normal through ghostscript.